# The jittering jets explosion mechanism (JJEM) in electron capture supernovae

### v1

@@ 3.2 The relevant convective zones
A note on the threshold CO core mass for electron capture supernovae is in place here. While some papers on the ECSN (e.g. Limongi et al. 2024) cite a value of $1.415M_{\odot}$, others claim to get an explosion with lower core masses, e.g., $1.34M_{\odot}$ (Stockinger et al., 2020b). The value is also highly dependent on the mixing properties (e.g. Zha et al. 2019 who also state a CO core value of $1.36M_{\odot}$ leading to an explosion). Our M88 model has a CO core mass of $M^{\rm CO}_{\rm core}=1.375$, which is inside the range of the values mentioned above.
@@
From Figure 6 we learn two interesting things. The first is that the accreted material that launches the jets originates in a helium layer, as studied by Gilkis & Soker (2016) for FeCCSNe, and not from the hydrogen-rich envelope as studied by, e.g., Antoni & Quataert (2022) and Antoni & Quataert (2023), for FeCCSNe. Namely, there is no ‘failed supernova’ here that is exploded only by the outer envelope layers. In any case, according to the JJEM there are no ‘failed supernovae’ (because jets explode all core-collapsing massive stars), a claim supported by recent observations (Byrne & Fraser, 2022; Strotjohann et al., 2023) and theoretical considerations (see review by Soker 2024).
@@ 5 Other models
${}^{16}$ 0.0911 0.15 0.112 0.2705 0.178
${}^{20}$e 0.061 0.0561 0.0665 0.1939 0.1162
${}^{24}$g 0.0081 0.01 0.01 0.0259 0.0171
${}^{28}$i$+^{32,34}$ 0.5245 0.5861 0.6004 0.3715 0.538
${}^{56}$e 0.06503 $3.9\times 10^{-4}$ $5.9\times 10^{-4}$ $3.6\times 10^{-6}$ $8.0\times 10^{-7}$
@@
There are two prominent outcomes from Table 1 that are relevant to our study in the frame of the JJEM. The first is that the final NS mass of an ECSN in all cases according to the criteria we use falls in the range of $1.25M_{\odot}<M_{\rm NS}<1.43M_{\odot}$. The very low density above the pre-collapse core ensures that even if the exploding layer is further out in the core the final NS mass will be $\simeq 1.4M_{\odot}$, compatible with observations. We note that Stockinger et al. (2020a) who simulated neutrino-driven explosion found the final NS mass in their ECSN models to be in the range of $1.195-1.23M_{\odot}$. The range we find here for the final NS masses might change following more accurate simulations and the inclusion of pre-collapse core rotation. However, we expect that the final NS masses that the JJEM predicts will stay somewhat larger than what the neutrino-driven explosion mechanism predicts.
@@ 6 Summary
Studies in recent years show that many, and probably most if not all, CCSNe are exploded by jets (Soker 2024 for a very recent review and references therein). The jet-driven explosion mechanism that can work in most CCSNe is the JJEM. Earlier studies of the JJEM concentrate on FeCCSNe (section 1). We, therefore, set the goal of examining whether the progenitors of ECSNe also have the convective layers that can supply the stochastic angular momentum to the intermittent accretion disks around the newly born NS that launch the exploding jets. To that goal, we simulated five stellar models that we summarized in Table 1.
For a pre-collapse convective layer, the ‘exploding layer’, to supply the stochastic angular momentum and mass of the intermittent accretion disks in CCSNe, we demand a minimum specific angular momentum parameter $j>j_{\rm jje}$ (equation 1), and minimum mass $M_{\rm acc}$ according to equation (2). We examined two values of $j_{\rm jje}=2.5\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$ and $j_{\rm jje}=20\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$; we expect the exploding layer to be somewhere in between these two. We found that the exploding layer resides in the helium-rich shell above the CO core (Figures 3 and 6). This is different than in FeCCSNe where the exploding layers are near the silicon-burning shell or the oxygen-burning shell of the pre-collapse core (e.g., Shishkin & Soker 2021). We found that the progenitors of ECSNe have convective layers in their helium-rich zone that fulfill the demands of the JJEM to set an explosion when they are accreted by the newly born NS.
We find that the exploding layer resides at a distance of $r_{\rm b}\approx 0.05-3R_{\odot}$ (Figures 3 and 6, and Table 1). This implies a free-fall time to the center of $f_{\rm ff,b}\approx{\rm few~{}10~{}sec}-{\rm few~{}hour}$. Because our simulations stop for numerical reasons before the core collapse phase, the actual radius of the exploding layers at the core collapse might be smaller. Nonetheless, the explosion phase might occur minutes to hours after core collapse, but still much shorter than the time the shock wave crosses the giant star. Therefore, the CCSN lightcurve is expected to be similar to that where the delay is only a few seconds. Other effects should be the topic of future simulations.
Overall, our results that ECSN progenitors have convective layers that can serve as the exploding layers of the JJEM even at the lower mass range of CCSNe (namely, ECSNe), add to the growing support (Soker, 2024) of the JJEM as the main explosion mechanism of massive stars.
@@ Acknowledgments
This research was supported by a grant from the Israel Science Foundation (769/20).

### v2

@@ 4.2 The relevant convective zones
A note on the threshold CO core mass for electron capture supernovae is in place here. While some papers on the ECSN (e.g. Limongi et al. 2024) cite a value of $1.415M_{\odot}$, others claim to get an explosion with lower core masses, e.g., $1.34M_{\odot}$ (Stockinger et al., 2020). The value is also highly dependent on the mixing properties (e.g. Zha et al. 2019 who also state a CO core value of $1.36M_{\odot}$ leading to an explosion). Our M88 model has a CO core mass of $M^{\rm CO}_{\rm core}=1.375$, which is inside the range of the values mentioned above.
@@
From Figure 6 we learn two interesting things. The first is that the accreted material that launches the jets originates in a helium layer, as studied by Gilkis & Soker (2016) for FeCCSNe, and not from the hydrogen-rich envelope as studied by, e.g., Antoni & Quataert (2022) and Antoni & Quataert (2023), for FeCCSNe. Namely, there is no ‘failed supernova’ here that is exploded only by the outer envelope layers. In any case, according to the JJEM there are no ‘failed supernovae’ (because jets explode all core-collapsing massive stars), a claim supported by recent observations (Byrne & Fraser, 2022; Strotjohann et al., 2024) and theoretical considerations (see review by Soker 2024b).
@@
We end this section by noting that some recent results suggest that even FeCCSNe in the frame of the delayed neutrino explosion mechanism require several seconds of explosion process (e.g., Bollig et al. 2021; Burrows & Vartanyan 2021), rather than only one second, to reach observed explosion energies. Overall, the CCSN explosion process, in some cases, might be much longer than the canonical one-second explosion timescale (which is the dynamical timescale of the collapsing inner core). This, in turn, allows the accretion of core layers outside the silicon core. In any case, the final mass of the proto-NS in ECSN or FeCCSN simulations can extend the $\simeq 1.2-2~{}M_{\odot}$ range, implying the accretion of outer core layers in most models.
@@ 6 Other models
16O 0.0911 0.15 0.112 0.2705 0.178
20Ne 0.061 0.0561 0.0665 0.1939 0.1162
24Mg 0.0081 0.01 0.01 0.0259 0.0171
28Si$+^{32,34}$ 0.5245 0.5861 0.6004 0.3715 0.538
56Fe 0.06503 $3.9\times 10^{-4}$ $5.9\times 10^{-4}$ $3.6\times 10^{-6}$ $8.0\times 10^{-7}$
@@
There are two prominent outcomes from Table 1 that are relevant to our study in the frame of the JJEM. The first is that the final NS mass of an ECSN in all cases according to the criteria we use falls in the range of $1.25M_{\odot}<M_{\rm NS}<1.43M_{\odot}$. The very low density above the pre-collapse core ensures that even if the exploding layer is further out in the core the final NS mass will be $\simeq 1.4M_{\odot}$, compatible with observations. We note that Stockinger et al. (2020) and Radice et al. (2017) who simulated neutrino-driven explosion found the final NS mass in their ECSN models to be in the range of $1.195-1.23M_{\odot}$ and $1.188M_{\odot}\lesssim M_{\rm NS}<1.4M_{\odot}$, respectively. The range we find here for the final NS masses might change following more accurate simulations and the inclusion of pre-collapse core rotation. However, we expect that the final NS masses that the JJEM predicts will stay somewhat larger than what the neutrino-driven explosion mechanism predicts.
@@ 7 Summary
Studies in recent years show that many, and probably most if not all, CCSNe are exploded by jets (e.g., Soker 2024a and Soker 2024b for a very recent review and earlier references therein). The jet-driven explosion mechanism that can work in most CCSNe is the JJEM. Earlier studies of the JJEM concentrate on FeCCSNe (section 1). Therefore, we aim to examine whether the progenitors of ECSNe also have the convective layers that can supply the stochastic angular momentum to the intermittent accretion disks around the newly born NS that launch the exploding jets. To that goal, we simulated five stellar models summarized in Table 1.
For a pre-collapse convective layer, the ‘exploding layer’, to supply the stochastic angular momentum and mass of the intermittent accretion disks in CCSNe, we demand a minimum specific angular momentum parameter $j>j_{\rm jje}$ (equation 1), and minimum mass $M_{\rm acc}$ according to equation (2). We examined two values of $j_{\rm jje}=2.5\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$ and $j_{\rm jje}=20\times 10^{15}{~{}\rm cm}^{2}{~{}\rm s}^{-1}$; we expect the exploding layer to be somewhere in between these two. This threshold value is a conjecture of the JJEM used by studies of the JJEM in FeCCSNe (see section 2), which we adopted here for ECSNe. It has yet to be validated by numerical simulations. We hope the new findings of point-symmetric CCSN remnants will motivate such studies.
We found that the exploding layer resides in the helium-rich shell above the CO core (Figures 3 and 6). This is different than in FeCCSNe where the exploding layers are near the silicon-burning shell or the oxygen-burning shell of the pre-collapse core (e.g., Shishkin & Soker 2021). We found that the progenitors of ECSNe have convective layers in their helium-rich zone that fulfill the demands of the JJEM to set an explosion when the newly born NS accretes them.
We find that the exploding layer resides at a distance of $r_{\rm b}\approx 0.05-3R_{\odot}$ (Figures 3 and 6, and Table 1). This implies a free-fall time to the center of $f_{\rm ff,b}\approx{\rm few\times~{}10~{}sec}-{\rm few\times~{}hour}$. Because our simulations stop for numerical reasons before the core collapse phase, the actual radius of the exploding layers at the core collapse might be smaller. Nonetheless, the explosion phase might occur minutes to hours after core collapse, but still much shorter than the time the shock wave crosses the giant star. Therefore, the CCSN lightcurve is expected to be similar to that where the delay is only a few seconds. Other effects should be the topic of future simulations.
Overall, our results that ECSN progenitors have convective layers that can serve as the exploding layers of the JJEM even at the lower mass range of CCSNe (namely, ECSNe) add to the growing support (Soker, 2024b) of the JJEM as the main explosion mechanism of massive stars.
@@ Acknowledgments
We thank an anonymous referee for detailed and helpful comments and suggestions. This research was supported by a grant from the Israel Science Foundation (769/20).